if coordinates of orthocenter and centroid are (4,-1)and (2,1) for atriangle, coordinates of the circumcentre are

Grade:11

1 Answers

pothamsetty chethan
6 years ago
We have the relation O=3G-2S(4,-1)=(6,3)-2S(-2,-4)=-2SS=(1,2)(Where the O is orthocenterS is circumventedG is centroid of the triangle